find the equation that passes through the point (2,1) and perpendicular to the line y=5x-3

Answer by mananth(16949)

1 y = 5 x -3
Divide by 1
y = 5 x -3
Compare this equation with y=mx+b
slope m = 5

The slope of a line perpendicular to the above line will be the negative reciprocal - 1/5
m1*m2=-1


m= - 1/5 ,point ( 2 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = -2/5 + b
b= 7/5
m= - 1/5
The required equation is y = - 1/ 5 x + 7/ 5
m.ananth@hotmail.ca